A solid concrete wall 4.0 m by 2.4 m and 30 cm thick, with a thermal conductivity of 1.3 W/(m
Answer #1
A=4*2.4=9.6 m^2
t=0.30 m
dT=18-6=12
q=kAdT/dx=kA*dT/t
q=1.3(9.6)(12)/0.30=4992 W=499.2 J/s
In 1 hour or 3600 sec
P=q*T=(499.2 J/s)(3600 s)=1797120 J=1.8 MJ
Answer #2
A=4*2.4=9.6 m^2
t=0.30 m
dT=18-6=12
q=kAdT/dx=kA*dT/t
q=1.3(9.6)(12)/0.30=4992 W=499.2 J/s
In 1 hour or 3600 sec
P=q*T=(499.2 J/s)(3600 s)=1797120 J=1.8 MJ
Answer #3
check it
http://www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=10&ved=0CHYQFjAJ&url=http%3A%2F%2Fjosefo-ciencia.com%2F4o%2Fanschp18.doc&ei=EHe6UIXgPMWUjAK4_4HIBw&usg=AFQjCNGOePtHrzBeAK3BYcuQ6EjBePopZQ&sig2=HAFYnSLvEp7Se3LY8fd8_w